A rubber ball of mass 50g falls from a height of 5m and rebounds to a height of 1.25m. Find the impluse and the average force between the ball and the ground if the time for which they are in contact was 0.1sec?

Grade:12th pass

1 Answers

Khimraj
5 years ago
velocity of ball before it comes in contact with ground
v1 = \sqrt{2*g*5} = 10m/s
velocity of ball after it comes in contact with ground
v2 = \sqrt{2*g*1.25} = 5m/s
so change in velocity = 10 – (-5) = 15m/s
So Impuse = m\Deltav = 0.05*15 = 0.75 Kg*m/s
and average Force = Impluse/time = 0.75/0.1 = 7.5N
Hope it clears. If you like answer then please approve it.